Find the value of b for which the equation has exactly one real root. (Hint- think discriminate0

x^2+bx+36=0

Answer by scott8148(6628)

if b is twice the square root of the constant term, then the quadratic is a perfect square (one real root)

b=2sqrt(36) __ b=2(+- 6) __ b=+- 12


if the discriminant equals zero, then there is one real root

b^2-4ac=0 __ b^2=4ac __ b=+- sqrt(4ac) __ b=+- sqrt(4*1*36) __ b=+- sqrt(144) __ b=+- 12